if in a GP , R denotes the sum of reciprocals of terms , S denotes  sum of the terms and P denotes the product of the terms . Then what is P2R3:S= ?

Grade:11

1 Answers

moumi roy
7 years ago
ANS=1
LET TERMS BE a/r,a,ar
R=(r/a)+(1/a)+(1/ar)=(r2+r+1)/ar
S=(a/r)+a+ar=a(r2+r+1)/r
P=a3
putting in (P^2*R^3)/s^3 we get 1